OpenStudy (anonymous):

If a line crosses the y-axis at (0, 1) and has a slope of 45, what is the equation of the line?

OpenStudy (anonymous):

4y − 5x = 5 y − 4x = 5 5y + 4x = 5 5y − 4x = 5

OpenStudy (anonymous):

Do you know?

OpenStudy (igreen):

Um...hold on.

OpenStudy (igreen):

D